find the equation of a line that passes through 4,8 and -3-6 my problem is when she asked questions like this one how do you now which formula to use thank you
Answer by CharlesG2(828)

find the equation of a line that passes through 4,8 and -3-6 my problem is when she asked questions like this one how do you now which formula to use thank you
the equation of a line is y=mx+b, where m is the slope, and b is the y-intercept (sometimes called the verical intercept), b is the point (0,b)
m=(y2-y1)/(x2-x1) or the rise/run
am hoping -3-6 means (-3,-6)
y=((-6-8)/(-3-4))x+b
y=(-14/-7)x+b
y=2x+b
2=(y2-y1)/(x2-x1)=(8-y1)/(4-0) (solving for y1 which will be b)
2=(8-y1)/4 ; 2=(-6-y1)/(-3-0)
8=8-y1 ; 2=(-6-y1)/-3
0=-y1 ; -6=-6-y1
0=y1 ; 0=-y1-->0=y1 -----> b is 0
y=2x